# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing

### v3

@@ 1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, regardless of public or private transactions it serves. As shown in recent studies [28], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the DApps relying on real-time blockchain access.
@@
This work takes a systematic and semi-automated approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is a promising approach but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool, which resides in the pre-consensus phase, needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees or prices are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [36], Loki [29] and Tyr [25] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs as evaluated in Appendix § A, let alone discover more sophisticated new ADAMS bugs.
@@
We propose symbolized-stateful mempool fuzzing or mpfuzz. To begin with, mpfuzz is set up and run in a three-step workflow: It is first manually set up against a size-reduced mempool under test, which induces a much smaller search space for fuzzing. Second, mpfuzz is iteratively run against the reduced mempool to discover short exploits. Third, the short exploits are manually extended to actual ones that are functional on real mempools and Ethereum clients. Internally, the design of mpfuzz is based on seven transaction symbols we design out of Ethereum semantics: future, parent, overdraft, latent overdraft, and replacement transactions. Under these symbols, in each iteration of fuzzing, mpfuzz explores one concrete transaction per symbol, sends the generated transaction sequence to the target mempool, observes the mempool end state, and extracts the feedback of symbolized state coverage and state promisingness in reaching bug oracles to guide the next round of fuzzing. In particular, mpfuzz employs a novel technique to evaluate state promisingness, that is, by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
@@
Found attacks: We run mpfuzz on six leading execution-layer Ethereum clients deployed on the mainnet’s public-transaction path (Geth [7], Nethermind [11], Erigon [5], Besu [10], Reth [13], and OpenEthereum [12]), three PBS builders (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot $v1.11.5$ [6], EigenPhi [4] and bloXroute [1]), and three Ethereum-like clients (BSC $v1.3.8$ [2] deployed on Binance Smart Chain, go-opera $v1.1.3$ [9] on Fantom, and core-geth $v1.12.19$ [3] on Ethereum Classic). Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude. On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
@@
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc. We propose rule-based mitigation schemes against all newly discovered vulnerabilities.
@@ 2 Related Works
Loki [29] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. That is, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and GasPrice are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
@@ 3 Background
The core mempool design is transaction admission: Given an initial state, whether and how the mempool admits an arriving transaction. In practice, example policies include those favoring admitting transactions of higher prices (and thus evicting or declining the ones of lower prices), transactions arriving earlier, certain transaction types (e.g., parent over child transactions), and valid transactions (over future transactions). See different types of Ethereum transactions in § 4.1.
@@ 4 Threat Model and Bug Oracle
The attacker’s goal is to disable the mempools on the critical or all nodes in an Etheurem network and cause damage such as dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services. In practice, this damage is beneficial to competing block validators, MEV searchers, and RPC services. Besides, the attacker aims to cause this damage at asymmetrically low costs, as will be described.
4.1 Notations
Transactions: In Ethereum, we consider two types of accounts: a benign account of index $i$ denoted by $B_{i}$, and an adversarial account of index $i$ denoted by $A_{i}$. A (concrete) Ethereum transaction $tx$ is denoted by $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$}}$, where the sender is $s$, the nonce is $n$, the GasPrice is $f$, and the transferred value in Ether is $v$. This work does not consider the “data” field in Ethereum transactions or smart contracts because smart contracts do not affect the validity of the Ethereum transaction. The notations are summarized in Table 1.
Table 1: Notations
@@ 4.1 Notations
4.2 Exploit Discovery Workflow
To set up the stage, we present the workflow overview in this work. The workflow is depicted in Figure 1. To discover vulnerabilities and exploits in an actual mempool, one has to 0) set up mpfuzz by various parameters (as will be introduced), 1) reduce the mempool from its original configuration to a much smaller version, named mempool under test (MUT). 2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 7.1). Fuzzing leads to the discovery of short exploits. 3) Short exploits are extended to the longer ones, that is, actual exploits. 4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 6.3.1) or a network we set up locally running unmodified Ethereum clients (see § 4 and § D). Fuzzing a reduced mempool (MUT) instead of an original one is necessary to ensure the efficient execution of the mpfuzz, which entails many iterations of re-initiating and populating the mempool. We discuss and evaluate the possible false positives that could be introduced by fuzzing smaller MUT in § 7.2.
In this workflow, fully automated is Step 2), that is, the discovery of short exploits on a given MUT (colored green in Figure 1). Other steps colored in gray require manual efforts and are described below: In Step 0), mpfuzz setup entails setting parameters in bug oracles (e.g., $\epsilon$ and $\lambda$) and fixed symbols in guiding fuzzing. The setup is usually one-time. In Step 1), mempool reduction entails reconfiguring the mempool to a much smaller capacity (e.g., MUT length $m=6$ or $m=16$, compared to the original size like $m^{\prime}=6144$ as in Geth). In addition, policy-specific parameters in the actual mempool are tuned down proportionally. For instance, the Geth mempool buffers up to $py_{1}^{\prime}=1024$ future transactions, and when buffering more than $py^{\prime}_{3}=5120$ pending transactions, the mempool triggers a protective policy to limit the number of transactions sent from the same sender by $py_{2}^{\prime}=16$ In a reduced MUT of $m=6$, these parameters are reduced while retaining the same proportion, such as $py_{1}=1,py_{3}=5,py_{2}=2$. The mempool reduction requires expert understanding and manual efforts, and it is by nature the best effort. In Step 3), it first de-duplicates the raw short exploits found by mpfuzz. Then, for each distinct short exploit, it extends it by repeating the transaction-admission events to fit into the actual mempool, during which one may practice strategies like increasing nonce, switching sender accounts, and measuring normal transactions in the actual mempool. In exploit extension, one could also consider the factors not modeled in MUT or mpfuzz such as $BasePrice$, as seen in the case in § 6.4.
4.3 Bug Oracle: Definitions & Rationale
@@
The transaction admission timeline is simplified from the timeline that an operational mempool experiences where transaction arrival and block building can be interleaved. The simplification is intended to trade off the accuracy of fuzzing results for fuzzing efficiency as discussed in § 7.2.
@@
Particularly, in the threat model of mempool DoS, we assume the validators are benign and functional in building blocks. Because in all Ethereum clients we know, transactions are admitted into mempool without execution (i.e., executing the smart contracts they invoke), GasUsed is not a factor in transaction admission. [^2] The further design rationale for admitting transactions without execution is that an Ethereum transaction $tx$’s GasUsed is non-deterministic unless the ordering of $tx$ w.r.t. other transactions in the block is fixed and $tx$ is executed. Thus, our bug oracle does not vary GasUsed. Instead, all the transactions are fixed at $21000$ Gas (i.e., as if they don’t invoke smart contracts). [^3] We may use price and fees interchangeably. This design significantly reduces the search space without loss in covering different mempool behaviors. Besides, our bug oracle does not model the effect of EIP1559 or $BasePrice$. Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $PriorityPrice$ or damage on validator revenue.
@@ 5.1 Transaction Symbolization
Table 2: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
@@
Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed GasPrice $3$ wei and of value $1$ Ether. That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table 2. Among all symbols, $\mathcal{N}$ is the only symbol associated with benign sender accounts.
@@
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$. The instantiated child transaction is fixed at GasPrice of $m+4$
@@
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 2, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$). The design rationale of transaction symbolization is deferred to Appendix A.1.
5.1.1 State Search Algorithm

### v4

# Towards Automated Discovery of Asymmetric Mempool DoS in Blockchains

@@ 1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, regardless of public or private transactions it serves. As shown in recent studies [1], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
@@
This work takes a systematic and semi-automated approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is a promising approach but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool, which resides in the pre-consensus phase, needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees or prices are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [7], Loki [8] and Tyr [9] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs, let alone discover more sophisticated new ADAMS bugs.
@@
We propose symbolized-stateful mempool fuzzing or mpfuzz. To begin with, mpfuzz is set up and run in a three-step workflow: It is first manually set up against a size-reduced mempool under test, which induces a much smaller search space for fuzzing. Second, mpfuzz is iteratively run against the reduced mempool to discover short exploits. Third, mpfuzz automatically extends the short exploits to actual ones that are functional on real mempools and Ethereum clients. Internally, the design of mpfuzz is based on seven transaction symbols we design out of Ethereum semantics: future, parent, overdraft, latent overdraft, and replacement transactions. Under these symbols, in each iteration of fuzzing, mpfuzz explores one concrete transaction per symbol, sends the generated transaction sequence to the target mempool, observes the mempool end state, and extracts the feedback of symbolized state coverage and state promisingness in reaching bug oracles to guide the next round of fuzzing. In particular, mpfuzz employs a novel technique to evaluate state promisingness, that is, by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future. After finding a short exploit against the reduced mempool, mpfuzz automatically extends the small exploit to a more extensive exploit that is effective within a larger mempool. Specifically, mpfuzz extracts all the unique admission event patterns in the small exploit and aligns these patterns to construct the attack transaction sequence on the larger-size mempool.
@@
Found attacks: We run mpfuzz on six leading execution-layer Ethereum clients deployed on the mainnet’s public-transaction path (Geth [10], Nethermind [11], Erigon [12], Besu [13], Reth [14], and OpenEthereum [15]), three PBS builders (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot $v1.11.5$ [16], EigenPhi [17] and bloXroute [18]), and three Ethereum-like clients (BSC $v1.3.8$ [19] deployed on Binance Smart Chain, go-opera $v1.1.3$ [20] on Fantom, and core-geth $v1.12.19$ [21] on Ethereum Classic). On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
@@
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
@@
Systematic evaluation: We conducted an ablation study to evaluate the performance of mpfuzz, which incorporates three distinct techniques. We aim to identify the contribution and significance of each component to the overall mpfuzz. We implement four baseline fuzzers to compare against mpfuzz. The first baseline is a stateless fuzzer built within the GoFuzz framework [24]. For the remaining three baselines, each one selectively disables a single technique from mpfuzz: symbolized state, symbolized input, and promising feedback and energy. The experimental result shows that the stateless fuzzer is the least effective in discovering ADAMS exploits. Among the three distinct techniques, the symbolized input technique improves efficiency more compared to that of the promising feedback and symbolized state coverage, and the state promisingness in feedback is more beneficial for performance than symbolized state coverage.
We systematically evaluate the attack success rate and cost of all the exploits found by mpfuzz on the Ethereum testnet and a local network. Specifically, we evaluate a newly discovered eviction attack on Goerli testnet. The result of the eviction attack shows the Gas used by normal transactions significantly drops to $3.19\%$ after the launch of the attack, which indicates that only a few normal transactions are included in the blocks following the attack. The locking attack evaluated on Goerli testnet shows the normal transactions received from the victim node drop essentially to zero, which indicates the locking attack is highly successful. Additionally, we evaluate other exploits on a local network. The evaluation results of all attacks on the Ethereum clients, PBS and Ethereum-like clients show that the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block.
@@
$\bullet$ New discovery of mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients, PBS and Ethereum-like clients. We found and reported $24$ newly discovered ADAMS bugs. After our reporting, $15$ bugs are confirmed and $3$ bugs are fixed.
$\bullet$ New found attack evaluation: This paper presents the evaluation results of the attack success rate and cost of all the exploits found by mpfuzz on the Ethereum testnet and a local network. By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
@@ 2 Related Works
Loki [8] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. In other words, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and Gas price are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
@@ 3 Background
The core mempool design is transaction admission: Given an initial state, whether and how the mempool admits an arriving transaction. In practice, example policies include those favoring admitting transactions of higher prices (and thus evicting or declining the ones of lower prices), transactions arriving earlier, certain transaction types (e.g., parent over child transactions), and valid transactions (over future transactions). See different types of Ethereum transactions in § 4-A.
@@ 4 Threat Model and Bug Oracle
The attacker’s goal is to disable the mempools on the critical nodes or all nodes in an Etheurem network and cause damage such as deployed or dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services like MEV searchers. In practice, this damage is beneficial to competing block validators, MEV searchers, and RPC services. Besides, the attacker aims to cause this damage at asymmetrically low costs, as will be described.
4-A Notations
Transactions: In Ethereum, we consider two types of accounts: a benign account of index $i$ denoted by $B_{i}$, and an adversarial account of index $i$ denoted by $A_{i}$. A (concrete) Ethereum transaction $tx$ is denoted by $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$}}$, where the sender is $s$, the nonce is $n$, the Gas price is $f$, and the transferred value in Ether is $v$. This work does not consider the “data” field in Ethereum transactions or smart contracts because smart contracts do not affect the validity of the Ethereum transaction. The notations are summarized in Table I.
TABLE I: Notations
@@ 4-A Notations
4-B Exploit Discovery Workflow
To set up the stage, we present the workflow overview in this work. The workflow is depicted in Figure 1. To discover vulnerabilities and exploits in an actual mempool, one has to 0) set up mpfuzz by various parameters (as will be introduced), 1) reduce the mempool from its original configuration to a much smaller version, named mempool under test (MUT). 2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 8-A). Fuzzing leads to the discovery of short exploits. 3) Short exploits are extended to the longer ones, that is, actual exploits. 4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 7-A1) or a network we set up locally running unmodified Ethereum clients (see § 7-B2, § 7-C and § 7-D ). Fuzzing a reduced mempool (MUT) instead of an original one is necessary to ensure the efficient execution of the mpfuzz, which entails many iterations of re-initiating and populating the mempool.
In this workflow, fully automated is Step 2) and Step 3), that is, the discovery of short exploits on a given MUT and extension of the short exploit to an actual exploit (colored green in Figure 1). Other steps colored in gray require manual efforts and are described below: In Step 0), mpfuzz setup entails setting parameters in bug oracles (e.g., $\epsilon$ and $\lambda$) and fixed symbols in guiding fuzzing. The setup is usually one-time. In Step 1), mempool reduction entails reconfiguring the mempool to a much smaller capacity (e.g., MUT length $m=6$ or $m=16$, compared to the original size like $m^{\prime}=6144$ as in Geth). In addition, policy-specific parameters in the actual mempool are tuned down proportionally. For instance, the Geth mempool buffers up to $py_{1}^{\prime}=1024$ future transactions, and when buffering more than $py^{\prime}_{3}=5120$ pending transactions, the mempool triggers a protective policy to limit the number of transactions sent from the same sender by $py_{2}^{\prime}=16$ In a reduced MUT of $m=6$, these parameters are reduced while retaining the same proportion, such as $py_{1}=1,py_{3}=5,py_{2}=2$. The mempool reduction requires expert understanding and manual efforts, and it is by nature the best effort.
4-C Bug Oracle: Definitions & Rationale
@@
The transaction admission timeline is simplified from the timeline that an operational mempool experiences where transaction arrival and block building can be interleaved.
@@
Particularly, in the threat model of mempool DoS, we assume the validators are benign and functional in building blocks. Because in all Ethereum clients we know, transactions are admitted into mempool without execution (i.e., executing the smart contracts they invoke), GasUsed is not a factor in transaction admission. [^2] The further design rationale for admitting transactions without execution is that an Ethereum transaction $tx$’s GasUsed is non-deterministic unless the ordering of $tx$ w.r.t. other transactions in the block is fixed and $tx$ is executed. Thus, our bug oracle does not vary GasUsed. Instead, all the transactions are fixed at $21000$ Gas (i.e., as if they don’t invoke smart contracts). [^3] We may use price and fees interchangeably. This design significantly reduces the search space without loss in covering different mempool behaviors. Besides, our bug oracle does not model the effect of EIP1559 or $BasePrice$. Without modeling $BasePrice$, our bug oracle can still capture the alternative transaction fee (the product of GasPrice and $21000$ Gas), which is still proportional to actual $tip$ or damage on validator revenue.
@@ 5-A Transaction Symbolization
TABLE II: Symbols, transactions, and associated costs. Tx refers to the instantiated transaction under a given symbol.
@@
Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed GasPrice $3$ wei and of value $1$ Ether. That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table II. Among all symbols, $\mathcal{N}$ is the only symbol associated with benign sender accounts.
@@
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+4}\end{bmatrix}$}}$. The instantiated child transaction is fixed at GasPrice of $m+4$.
@@
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 2, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+4}\end{bmatrix}$}}$).
5-A1 State Search Algorithm